The vendor has supplied a replacement switch. Engineers have attended the data centre and replaced the switch. All services have been tested and working as expected.


We've fixed the core issue, and are waiting for things to recover. VMs should now be available. Technicans identified the issue as a partially failed switch which did not trigger a full fail-over to the second switch. The affected switch was removed from path and internet connectivity has been restored to the affected services. The failed switch will be taken offline, logs reviewed and root caused investigated.

Technicians are attending the Perth data centre. We will provide an update at 8:15pm (WST).

This incident will also affect some Simtex calling as some of the VoIP servers are affected by this issue.


Selected virtual servers are experiencing network connectivity issues on some hosts preventing access to the internet. The team are currently investigating and will update as soon as possible.

Began at:

Affected components
  • Australia - Perth (Primary)
    • Compute
    • Simtex VoIP and Fax
    • Veeam Services
    • Zerto Services